# Written by AAAI Press Staff1

### v2

@@ 3.2 Action: Selection of Nodes and Partitions
Phase 1: Node selection to identify node to perturb
@@
$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in P^{t}-\operatorname{PART}(P^{t},v)}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=p|}}}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=\operatorname{PART}(P^{t},v)|}}$ (10)
Further, a node having a higher degree implies it has several edges associated to it, hence, an incorrect placement of it could contribute to higher partitioning value. Hence, we normalize the scores by the $degree$ of the node.
Phase 2: Inductive Method for Partition Selection
Once a node is selected, the next phase involves choosing the new partition for the node. Towards this, we design an approach empowered by Graph Neural Networks (GNNs) (hamilton2017inductive) which enables the model to be inductive with respect to size of graph. Further, instead of predicting a fixed-size score vector (nazi2019gap; tsitsulin2023graph) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too. We discuss both above points in section below.
@@
To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (hamilton2017inductive). GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{emb_{init}}(u)$ using eq. 8. We perform $L$ layers of message passing to compute representations of nodes. To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(hamilton2017inductive):
$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\frac{1}{|\mathcal{N}_{u}|}\sum_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (11)
@@
Scoring partitions. Recall from eq. 9, each partition at time $t$ is represented using the nodes belonging to that partition. Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$. In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap; tsitsulin2023graph), the proposed design choice makes the model inductive to the number of partitions. Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
@@
$\operatorname{PartScore}$ $(p,v){=}\operatorname{AGG}(\{\operatorname{MLP}(\sigma(h_{v}|h_{u}))$ (12)
@@
$\pi((a^{t}{=}p)/S^{t})=\frac{\exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (13)
During the course of trajectory of length $T$, we sample action $a^{t}\in\mathbb{R}$ i.e., the assignment of the partition for the node selected in phase 1 at step $t$ using policy $\pi$.
@@
$P^{t+1}_{i}\leftarrow P^{t}_{i}\setminus v\text{ and }P^{t+1}_{j}\leftarrow P^{t}_{j}\cup v$ (14)
3.3 Reward, Training & Inference
Reward. Our aim is to improve the value of the overall partitioning objective. One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$, i.e., $R^{t}=(Obj(\mathcal{G},\mathcal{P}^{t+1})-Obj(\mathcal{G},\mathcal{P}^{t}))/(Obj(\mathcal{G},\mathcal{P}^{t+1})+Obj(\mathcal{G},\mathcal{P}^{t}))$. Here, the denominator term ensures that the model receives a significant reward for even slight enhancements when dealing with a small objective value. However, this definition of reward focuses on short-term improvements instead of long-term. Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term (sutton2018reinforcement). The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term). Mathematically,
$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}$ (15)
@@ 3.3 Reward, Training & Inference
The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable. This is an advantage over existing neural methods (nazi2019gap; tsitsulin2023graph) where having a differentiable form of the partitioning objective is a pre-requisite.
Policy Loss Computation and Parameter Update. Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others. Towards this, we use REINFORCE gradient estimator (williams1992simple) to optimize the parameters of our policy network. Specifically, we wish to maximize the reward obtained for the horizon of length $T$ with discounted rewards $D^{t}$. Towards this end, we define a reward function $J(\pi_{\theta})$ as:
$J(\pi_{\theta})=\mathbb{E}\big{[}\sum_{t=0}^{T}\left(D^{t}\right)\big{]}$ (16)
We, then, optimize $J(\pi_{\theta})$ as follows:
${\nabla J(\pi_{\theta})=\left[\sum_{t=0}^{T}\left(D^{t}\right)\nabla_{\theta}log\pi_{\theta}({a}^{t}/S^{t})\right]}$ (17)
@@ 3.4 Time complexity
$\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$. Here $\alpha$ is the number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Further, for sparse graphs $\mathcal{E}=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$. For detailed derivation please see appendix  A.4.
4 Experiments
@@
  • Inductivity: As one of the major strengths, unlike existing neural models such as GAP (nazi2019gap), DMon (tsitsulin2023graph) and MinCutPool (bianchi2020spectral), our proposed method NeuroCUT is inductive on the number of partitions and consequently can generalize to unseen number of partitions.
4.1 Experimental Setup
Datasets:
We use four real datasets for our experiments. They are described below and their statistics are present in the appendix (please see Table 7).
  • Cora and Citeseer (sen2008collective): These are citation networks where nodes correspond to individual papers and edges represent citations between papers. The node features are extracted using a bag-of-words approach applied to paper abstracts.
  • Harbin (li2019learning): This is a road network extracted from Harbin city, China. The nodes correspond to road intersections and node features represent latitude and longitude of a road intersection.
  • Actor (platonov2023critical): This dataset is based upon Wikipedia data where each node in the graph corresponds to an actor, and the edge between two nodes denotes co-occurrence on the same Wikipedia page. Node features correspond to keywords in Wikipedia pages.
Partitioning objectives:
@@ 4.1 Experimental Setup
Neural baselines: We compare with DMon (tsitsulin2023graph), GAP (nazi2019gap), MinCutPool  (bianchi2020spectral) and Ortho (tsitsulin2023graph). DMon is the state-of-the-art neural attributed-graph clustering method. GAP is optimized for balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut. MinCutPool optimizes for normalized cut and uses an additional orthogonality regularizer. Ortho is the orthogonality regularizer described in DMon and MinCutPool.
Non-neural baselines: Following the settings of DMon (tsitsulin2023graph), we compare with K-means clustering applied on raw node features.
Other settings:
We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments. For NeuroCUT we used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$. We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$. For computing discounted reward in RL, we use discount factor $\gamma=0.99$. We set the length of trajectory $T$ during training as $2$. At time step $t$, the rewards are computed from time $t$ to $t+T$ and parameters of the policy $\pi_{\theta}$ are updated. The default number of anchor nodes for computing positional embeddings is set to $35$.
@@
Table 2: Results on Normalized Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
@@
Table 3: Results on Sparsest Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
4.2 Results on Transductive Setting
In the transductive setting, we compare our proposed method NeuroCUT against the mentioned baselines, where the neural models are trained and tested with the same number of partitions. Tables 2-5 present the results for different partitioning objectives across all datasets and methods. For the objectives under consideration, a smaller value depicts better performance. NeuroCUT demonstrates superior performance compared to both neural and non-neural baselines across four distinct partitioning objectives, highlighting its robustness. Further, we would also like to point out that in many cases, existing baselines produce “nan” values which are represented by “-” in the result tables.
Our method NeuroCUT incorporates dependency during inference and this leads to robust performance. Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in methods such as GAP (nazi2019gap). Further, unlike other methods, NeuroCUT also incorporates positional information in the form of Lipschitz embedding to better contextualize global node positional information. We also observe that the non-neural method K-means fails to perform on all objective functions since its objective is not aligned with the main objectives under consideration.
@@ 4.2 Results on Transductive Setting
Table 4: Results on k-MinCut. In most of the cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$. Values less than $10^{-2}$ are approximated to 0.
@@
Table 5: Results on Balanced Cut. In most cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
4.3 Results on Inductivity to Partition Count

### v3

@@ 3.2. Action: Selection of Nodes and Partitions
3.2.1. Phase 1: Node Selection to Identify Node to Perturb
@@
(10) $\operatorname{NodeScore}^{t}[v]=\\ \frac{\max_{p\in{P}^{t}\setminus\operatorname{PART}({P}^{t},v)}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}({P}^{t},u)=p|}}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}({P}^{t},u)=\operatorname{PART}({P}^{t},v)|}\times\frac{1}{degree(v)}$
For a node of interest $v$, the numerator computes the maximum number of neighbors in a different partition than that of $v$. As described in Table  1, $\operatorname{PART}(P^{t},v)$ refers to the partition of $v$ at step $t$. The expression $P^{t}\setminus\operatorname{PART}(P^{t},v)$ computes all other partitions except the partition of $v$. The term $|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=p|$ computes the number of neighbors of $v$ in the partition $p$. The denominator computes the number of neighbors of node $v$ in the same partition as $v$, referred to as $\operatorname{PART}(P^{t},v)$. Further, a node having a higher degree implies that it has several edges associated to it. Hence, an incorrect placement of it could contribute to a higher partitioning value. Therefore, we normalize the scores by the $degree$ of the node.
3.2.2. Phase 2: Inductive Method for Partition Selection
Once a node is selected, the next phase involves choosing the new partition for the node. Towards this, we design an approach empowered by Graph Neural Networks (GNNs) (Hamilton et al., 2017) which enables the model to be inductive with respect to size of graph. Further, instead of predicting a fixed-size score vector (Nazi et al., 2019; Tsitsulin et al., 2023b) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too. We discuss both above points in section below.
@@ 3.2.2. Phase 2: Inductive Method for Partition Selection
To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (Hamilton et al., 2017). GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{emb_{init}}(u)$ using eq. 8. We perform $L$ layers of message passing to compute representations of nodes. To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(Hamilton et al., 2017):
[…]
@@
Scoring Partitions. Recall from eq. 9, each partition at time $t$ is represented using the nodes belonging to that partition. Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$. In contrast to predicting a fixed-size score vector corresponding to number of partitions (Nazi et al., 2019; Tsitsulin et al., 2023b), the proposed design choice makes the model inductive to the number of partitions. Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
@@
(12) $\operatorname{PartScore}$ $(p,v){=}\operatorname{AGG}(\{\operatorname{MLP}(\sigma(h_{v}|h_{u}))$
@@
(13) $\pi((a^{t}{=}p)/S^{t})=\frac{\exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$
During the course of trajectory of length $T$, we sample action $a^{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathcal{P}}$ i.e., the assignment of the partition for the node selected in phase 1 at step $t$ using policy $\pi$.
@@
(14) $\hskip 43.36243ptP^{t+1}_{i}\leftarrow P^{t}_{i}\setminus\{v\}\text{ and }P^{t+1}_{j}\leftarrow P^{t}_{j}\cup\{v\}\vspace{-0.2in}$
3.3. Reward, Training & Inference
Reward. Our aim is to improve the value of the overall partitioning objective. One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$. Specifically,
(15) $R^{t}=\frac{(Obj(\mathcal{G},\mathcal{P}^{t})-Obj(\mathcal{G},\mathcal{P}^{t+1}))}{(Obj(\mathcal{G},\mathcal{P}^{t})+Obj(\mathcal{G},\mathcal{P}^{t+1}))}\cdot\lambda$
Here $\lambda$ is a hyperparameter that is used to scale the reward. The above reward expression incentivizes significant improvements in the objective function. This is achieved mathematically by considering both the change and the current value of the objective function. This design steers the model towards prioritizing substantial improvements, especially in low objective function regions, ultimately guiding it towards the overall minimum.
However, above definition of reward focuses on short-term improvements instead of long-term. Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term (Sutton and Barto, 2018). The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term). Mathematically,
(16) $D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}$
@@ 3.3. Reward, Training & Inference
The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable. This is an advantage over existing neural methods (Nazi et al., 2019; Tsitsulin et al., 2023b) where having a differentiable form of the partitioning objective is a pre-requisite.
Policy Loss Computation and Parameter Update. Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others. Towards this, we use REINFORCE gradient estimator (Williams, 1992) to optimize the parameters of our policy network. Specifically, we wish to maximize the reward obtained for the horizon of length $T$ with discounted rewards $D^{t}$. Towards this end, we define a reward function $J(\pi_{\theta})$ as:
(17) $J(\pi_{\theta})=\mathbb{E}\big{[}\sum_{t=0}^{T}\left(D^{t}\right)\big{]}$
We, then, optimize $J(\pi_{\theta})$ via policy gradient (Sutton and Barto, 2018) as follows:
(18) ${\nabla J(\pi_{\theta})=\left[\sum_{t=0}^{T}\left(D^{t}\right)\nabla_{\theta}log\pi_{\theta}({a}^{t}/S^{t})\right]}$
@@ 3.4. Time Complexity
$\mathcal{O}\big{(}(\alpha\times\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times|\mathcal{E}|})+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$. Here $\alpha$ is the number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$. For detailed derivation please see Appendix  A.2.
4. Experiments
@@
  • Inductivity: As one of the major strengths, unlike existing neural models such as GAP (Nazi et al., 2019), DMon (Tsitsulin et al., 2023b) and MinCutPool (Bianchi et al., 2020), our method NeuroCUT is inductive on the number of partitions and consequently can generalize to unseen number of partitions.
Code: Our code base is accessible at https://github.com/idea-iitd/NeuroCut.
4.1. Experimental Setup
4.1.1. Datasets:
We use four real datasets for our experiments. They are described below and their statistics are present in Table 2. For all datasets we use their largest connected component.
  • Cora and Citeseer (Sen et al., 2008): These are citation networks where nodes correspond to individual papers and edges represent citations between papers. The node features are extracted using a bag-of-words approach applied to paper abstracts.
  • Harbin (Li et al., 2019): This is a road network extracted from Harbin city, China. The nodes correspond to road intersections and node features represent latitude and longitude of a road intersection.
  • Actor (Platonov et al., 2023): This dataset is based upon Wikipedia data where each node in the graph corresponds to an actor, and the edge between two nodes denotes co-occurrence on the same Wikipedia page. Node features correspond to keywords in Wikipedia pages. This is a heterophilous dataset (Zheng et al., 2022) where nodes tend to connect to other nodes that are dissimilar.
  • Facebook (McAuley and Leskovec, 2012): In social network analysis, a $k$-way cut can delineate communities of users with minimal inter-community connections. This partitioning unveils hidden social circles and provides insights into how different groups interact within the larger network. The Facebook dataset, used for community detection, does not include node features. In this dataset, nodes represent users, and edges represent friendships.
  • Stochastic Block Models (SBM) (Abbe, 2017): SBMs are synthetic networks and produce graphs with communities, where subsets of nodes are characterized by specific edge densities within the community. This allows us to validate our method by comparing the identified partitions with the ground truth communities. This dataset does not include node features.
    Table 2. Dataset Statistics
    Dataset $|\mathcal{V}|$ $|\mathcal{E}|$Features Average Degree Clustering Coefficient Degree Assortativity
    Cora 2485 5069 1433 4.07 0.23 -0.07
    CiteSeer 2120 3679 3703 3.47 0.169 0.0075
    Harbin 6598 10492 2 3.18 0.036 0.22
    Actor 6198 14879 931 4.82 0.05 -0.048
    Facebook 1034 26749 - 51.7 0.526 0.431
    SBM 500 5150 - 20.6 0.18 -0.0059
4.1.2. Partitioning objectives:
@@ 4.1.3. Baselines:
Neural baselines: We compare with DMon (Tsitsulin et al., 2023b), GAP (Nazi et al., 2019), MinCutPool  (Bianchi et al., 2020) and Ortho (Tsitsulin et al., 2023b). DMon is the state-of-the-art neural attributed-graph clustering method. GAP is optimized for balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut. MinCutPool optimizes for normalized cut and uses an additional orthogonality regularizer. Ortho is the orthogonality regularizer described in DMon and MinCutPool. We also compare with DRL (Gatti et al., 2022) which solves for Normalized cut at $k=2$ in Appendix A.5
Non-neural baselines: Following the settings of DMon (Tsitsulin et al., 2023b), we compare with K-means clustering applied on raw node features. We also compare with standard graph clustering methods hMetis (Karypis et al., 1997) and Spectral clustering (Ng et al., 2001) in App Sec. A.4.
4.1.4. Other settings:
We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments. For NeuroCUT we used GraphSage(Hamilton et al., 2017) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$. We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$. For computing discounted reward in RL, we use discount factor $\gamma=0.99$. We set the length of trajectory $T$ during training as $2$. At time step $t$, the rewards are computed from time $t$ to $t+T$ and parameters of the policy $\pi_{\theta}$ are updated. The default number of anchor nodes for computing positional embeddings is set to $35$. We set $\beta=100$ and $c=0.85$ for Eq. 6. We set scaling factor $\lambda=100$ in eq. 15.
Table 3. Results on Normalized Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
@@ 4.1.4. Other settings:
Table 4. Results on Sparsest Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
@@
4.2. Results on Transductive Setting
In the transductive setting, we compare our proposed method NeuroCUT against the mentioned baselines, where the neural models are trained and tested with the same number of partitions. Tables 3-6 present the results for different partitioning objectives across all datasets and methods. For the objectives under consideration, a smaller value depicts better performance. NeuroCUT demonstrates superior performance compared to both neural and non-neural baselines across four distinct partitioning objectives, highlighting its robustness. Further, we would also like to point out that in many cases, existing baselines produce “nan” values which are represented by “-” in the result tables. This is due to the reason that every partition is not assigned atleast one node which leads to the situation where denominator becomes 00 in normalized, balanced and sparsest cut objectives as defined in Sec. 2.
Our method NeuroCUT incorporates dependency during inference and this leads to robust performance. Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in methods such as GAP (Nazi et al., 2019). Further, unlike other methods, NeuroCUT also incorporates positional information in the form of Lipschitz embedding to better contextualize global node positional information. We also observe that the non-neural method K-means fails to perform on all objective functions since its objective is not aligned with the main objectives under consideration.
Table 5. Results on k-MinCut. In most of the cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$. Values less than $10^{-2}$ are approximated to 0.
@@ 4.2. Results on Transductive Setting
Table 6. Results on Balanced Cut. In most cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
@@
4.3. Results on Inductivity to Partition Count
Table 7. Inductivity to unseen partition count. Here we set the target number of partitions $k=10$. I stands for inductive and T for transductive setting. In this table, the transductive version of NeuroCUT trained on $k=10$ serves as a point of reference when assessing the quality of the inductive version of NeuroCUT.